### v2

# GroundGazer: Camera-based indoor localization of mobile robots with millimeter accuracy at low cost

Abstract
Highly accurate indoor localization systems with mm positioning accuracy are currently very expensive. They include range finders (such as LiDAR), tachymeters, and motion capture systems relying on multiple high-end cameras. In this work, we introduce a high-accuracy, planar indoor localization system named GroundGazer (GG) for autonomous mobile robots (AMRs). GG estimates the AMR’s position with mm and its heading with sub-degree accuracy. The system requires only a monocular (fisheye) camera, a chessboard floor, and an optional laser diode. Our system is simple and low-cost, easy to set up, portable, robust, scalable to large areas and robot swarms, and potentially extendable to 3D position and orientation estimation.
Keywords— Indoor localization, mobile robots, computer vision
I Introduction
High-accuracy planar indoor localization applications, e.g. docking in warehousing and benchmarking low-accuracy indoor positioning systems with up to cm localization accuracy, require localization approaches with reliable mm positioning accuracy. Currently, such systems are very costly (generally $1000\,\mathdollar$, with costs rising quickly with increasing size of the coverage area). In this work, we introduce a novel camera-based indoor localization system for AMRs which achieves mm localization and sub-degree heading estimation accuracy with high robustness against occlusion of the position references on the ground at multitudes lower cost than existing systems of similar precision. We name it GroundGazer, as the camera placed on the AMR gazes at the passive position references on the ground.
Before describing the GG system, we give an introduction into currently existing indoor localization methods and sensor technologies for AMRs with mm positioning accuracy. For an overview of general indoor localization technologies, we refer to the surveys in [review_indoor_pos_habil] and [review_indoor_loc]. Comprehensive overviews of optical indoor localization systems can be found in [Survey_Optical_pos] and [cv_loc_survey].
I-A Localization technologies with mm positioning accuracy
Motion capture systems consist of passive infrared (IR) markers placed on the to be tracked AMR. With one marker, 3D positioning is possible. A minimum of three non-collinear markers enable 3D position and orientation estimation. Motion capture systems like the commercial systems by Qualisys and Vicon [^1] Available at: https://www.qualisys.com/ and https://www.vicon.com/. are very popular in robotics and industry. While capable of sub-mm 3D localization and sub-degree orientation estimation, they require line of sight (LOS), multiple high-end IR cameras, and become very expensive if large areas are to be observed. Further, the IR cameras not installed around the AMR complicate the time synchronization when the motion capture system is used as reference for another localization system with sensors installed on the AMR. Highly accurate time synchronization is required, since otherwise the mm localization accuracy is lost through badly matched time stamps.
A different, less frequently used solution for 3D position estimation with mm or sub-mm accuracy are tachymeters, e.g. used as 3D position reference system in [esparagos]. Tachymeters enable long range sensing, but like motion capture systems, they are expensive and require LOS and more complex time synchronization mechanisms. Another, novel positioning system with long distance capabilities based on interferometry is developed by Koherent [^2] Available at: https://koherent.io/.
Radio frequency (RF)-based backscatter systems using frequency modulated continuous wave (FMCW) radar and reflective tags (e.g. Van-Atta arrays) that modulate the reflected radar signal can also achieve 3D mm localization accuracy using high-precision range estimates of the tags and multilateration (MLAT) [hawkeye] or triangulation for positioning [mm_siemens].
In acoustic ultrasound systems, small timing errors in the measured time of flight (ToF) of the signals or in the synchronization between anchors cause lower positioning errors than in RF based ultra wideband (UWB) systems. The reason lies in the much lower propagation speed of acoustic waves. The company ZeroKey claims real-time and up to 1.5 mm 3D positioning accuracy with their commercial Quantum RLTS 2.0 system, given 6 acoustic anchors with ideal geometric placement [zerokey]. Positioning is performed with multiple active acoustic anchors and ToF measurements between anchors and the to be tracked AMR and by applying MLAT for localization. Since no product prices are available on their website or data sheets, we believe it likely not being a low-cost solution for AMR localization.
One of the most widely used localization sensor is LiDAR [lidar]. While high-end LiDAR systems can achieve 3D mm accuracy, they are also expensive and need precise maps of the environment through costly measurement campaigns to achieve accurate, absolute position estimates. Further, their accuracy degrades with dynamic changes of the environment that are not contained in the previously recorded reference map.
Deep learning based channel state information (CSI) fingerprinting with a massive multiple input multiple output (MIMO) system has recently also shown mm accuracy in a small area and under LOS conditions [csi]. However, similar as with LiDAR, the performance of such system degrades with changes in the environment, causing domain shifts between the measured fingerprints during inference and the pre-recorded fingerprint database.
Fiducial markers pose another solution for highly precise localization. In [ODOP], the combination of high-end binocular cameras and custom markers achieves sub-mm localization accuracy. Downsides are again the high cost of the cameras and the required LOS. The authors of [QR_docking_submm] proposed a docking system with sub-mm 2D positioning accuracy using quick response (QR) codes and an RGB camera. Although low-cost, the very short camera-QR code distance and the need for many QR codes limit its use for general AMR localization.
Additional systems based on visual markers, laser projections, magnetic fields, laser trackers, pseudolites, and infrared laser beams (iGPS) with (sub-)mm accuracy can be found in [Survey_Optical_pos] and [mautz_alternative_systems].
I-B Contributions
We introduce a novel mm accuracy localization approach with the intended application to serve as reference for less accurate indoor positioning systems. Our GG prototype described in this paper requires only a chessboard floor (or similar grid pattern), a camera (here a global shutter RGB fisheye camera), and an optional laser diode, i.e. low-cost off-the-shelve components. Currently, it only allows for planar position ($x$, $y$) and heading/yaw $\theta$ estimation, which is sufficient for our application as reference system for planar AMR position and heading estimation with known or irrelevant height, roll and pitch. In Sec. VI, we discuss an adaption that yields 3D position and orientation estimates, but has higher computational complexity and possibly lower accuracy.
The idea for our GG system originates from [ground_truth_laser]. They developed a fiducial marker localization system that requires a chessboard floor in addition to the fiducial markers, where a car is equipped with multiple fiducial markers placed on a plate mounted at the front of the car. To evaluate the positioning accuracy of this system, they installed two point laser diodes that project two points from the front and the back of the car onto the chessboard floor beneath it. The corner coordinates of the chessboard serve as position references and they measure the two laser point locations on the floor as ground truth by hand. The laser point at the front of the car is used as vehicle reference position and the heading is derived from the orientation of the line connecting both points w.r.t. to the chessboard. Our GG system builds on that idea, resulting in a highly accurate, camera-based indoor localization system.
I-C System overview
Our localization system is depicted in Fig. 1(a).
((a)) Our system in the laboratory equipped with a chessboard floor and QR codes. The AMR has a ground facing camera and a laser diode that projects a crosshair on the floor at the front.
((b)) 2D position and heading estimation using the detected crosshair (P5) and chessboard square corners (P1-P4). The globally known square corner coordinates through the QR codes enable accurate, global positioning.
Figure 1: Overview of our GG localization system in a) and conceptual depiction of the localization method in b).
It consists of an AMR, which is equipped with an RGB fisheye camera ($100\,\mathdollar$) with large field-of-view (FoV) and global shutter. Global shutter cameras have lower motion blur than rolling shutter cameras due to parallel collection of all pixels. The camera is facing the ground and a green laser diode (negligible cost) projects a crosshair on said ground. The ground itself is made up of a chessboard pattern with well known grid size.
The localization system works by detecting the crosshair in a square of the chessboard. We specifically selected a blue and white chessboard floor and a green laser diode, as green lasers are most easily visible and other commercially available black and white chessboard floors complicate the crosshair detection, caused by the large difference in laser reflective behavior between black and white squares. The laser diode is installed such that the AMR moves in the direction of the vertical crosshair line, thus the angle between the vertical crosshair line and the chessboard lines in $x$ direction gives the AMR’s heading. The corner coordinates of the chessboard squares are known and serve as passive position references for AMR localization.
Since the chessboard corner coordinates and the crosshair orientation are ambiguous to the camera which only observes a small, local area of the periodic chessboard pattern, we additionally equip the floor with different QR codes as global position and orientation references. All QR codes are oriented in the same direction, parallel to the chessboard lines, and their corners can therefore be used to extract the $x$- and $y$-directions of the chessboard lines in global coordinates. The QR code IDs allow for identification of the QR code squares and global positioning. By measuring the distance in $x$- and $y$-direction between a detected QR code square center and the center of the square in which the crosshair lies (neglecting the trivial case where they are identical), we can further identify the crosshair square and its known corner coordinates without requiring a QR code in each square. Measuring the relative position of the crosshair w.r.t. its square corners (at least three corners are required to build a square coordinate system, but we use four corners as known from a detected square) gives the AMR’s position and the orientation of the vertical crosshair line w.r.t. the chessboard lines in $x$ direction its heading. The AMR position and heading is set as the crosshair position and orientation. An image describing the AMR positioning is shown in Fig. 1(b).
High position and heading accuracy is enabled by placing the camera close to the ground and by precise estimation of the crosshair center and the crosshair lines, and the square corners and the chessboard lines. For this prototype, a camera with fisheye lens was chosen because of their large FoV, allowing observation of multiple chessboard squares with a small camera-ground distance. The whole processing chain, which transforms the initial fisheye image, detects the QR codes, the crosshair, and the chessboard squares, and finally estimates and smooths the AMR’s position and heading, is described in Sec. III.
II Related works with ground facing cameras
In this section, we focus on closely related systems that employ ground facing cameras for global localization. A general overview of indoor positioning systems with mm accuracy for AMRs was given in Sec. I-A.
One approach to global localization using ground facing cameras is by avoiding external infrastructure, i.e. localization using only the given ground texture. Relative positioning can be realized through simultaneous localization and mapping (SLAM) [prob_robotics], but we are interested in absolute, global positioning. Such positioning is possible with feature-matching of images collected during inference with a pre-collected image database of the ground [high_precision_matching, streetmap, binary_matching, feature_matching_newest]. The advantage of this is the applicability in (ideally) any kind of environment without requiring new infrastructure, making it a practical method e.g. for warehousing. However, feature-matching requires an expensive measurement campaign for collection of reference images of the floor with accurate ground truth, e.g. from LiDAR or motion capture systems. These approaches have also shown varying reliability for different surfaces [high_precision_matching] and the position errors are in the low cm range (ignoring outliers, but also not yet using temporal correlations) and the heading errors in the low degree range [feature_matching_newest]. Position errors further seem to increase with coverage area (map size) [feature_matching_newest]. The commercial feature-matching based systems by Accerion are reported to deliver sub-mm accuracy, but only for repeated tasks such as docking, where the task was accurately executed before, while stating cm accuracy for global localization [^3] Due to the limited product specifications on the company website (no published data sheets), we are relying on the information given on their Youtube channel found at https://www.youtube.com/@accerion4564.. While feature-matching approaches require no additional infrastructure, we argue that using known position references is the more suitable approach for reference systems due to higher accuracy, robustness, easier setup, and lower hardware requirements.
Various approaches have considered placing infrastructure on the floor for localization. In [floor_aruco], ArUco markers [ArUco] are distributed on the floor and combined with odometry in an Extended Kalman filter (EKF) for localization. However, they don’t give localization accuracy results. QR codes are also commonly applied, e.g. for AMRs in [QR_amr]. In [QR_drones], a drone with a ground facing camera was localized with a QR code grid on the floor and using the nearest QR code for position and heading estimation, showing cm accuracy. Different fiducial marker types are benchmarked regarding their detection behavior in ground based localization in [fiducial_floor_det].
It should be noted, however, that we believe mm accuracy with floor installations of ArUco, QR codes, or similar fiducial markers is possible. With ceiling installations of ArUco markers and much larger camera-marker distances, we could already achieve low cm accuracy in previous work [icar25_fiducial] and [QR_docking_submm] reports sub-mm accuracy with very small camera-QR code distance. Similar commercial systems can also be found. One example is the company Hikrobot, which provides AMRs and vision based localization solutions using a visual marker grid on the floor for warehousing, logistics etc. [^4] Available at: https://www.hikrobotics.com/en/.
Opposed to existing systems, we can deliver absolute 2D position and heading estimates with mm and sub-degree accuracy at low cost and with minimal setup. Our processing can likely also be sped-up through further optimization as discussed in Sec. VI to enable near real-time mm accuracy during online processing with common AMR velocities of $1\text{\,}\frac{m}{s}$, factoring in additional delays from exposure time and image transmission to the processor. This online bottleneck might be solved by adding high-frequency odometry. Neither SLAM nor feature-matching can produce this due to higher processing times and delayed position estimates. To the best of our knowledge, no system like GG exists right now for indoor AMR localization that offers a similar combination of high accuracy, low cost, simplicity, easy set up, scalability, portability, and fast processing.
III GroundGazer
III-A Image transformation
The first step of our processing is the transformation of the distorted and skewed image collected by the fisheye camera to a top-down view of the chessboard floor. This step consists of two transformations. First, the fisheye image needs to be undistorted. We apply the built-in undistort function from OpenCV [opencv_library] [^5] If not specified otherwise, we always refer to OpenCV functions for the remainder of the paper.. This function takes as input the distorted fisheye image, the intrinsic matrix $\mathbf{K}$ of the camera, and its distortion coefficient vector $\underline{d}$, and outputs the undistorted image. The intrinsic matrix $\mathbf{K}$ and the distortion coefficient vector $\underline{d}$ have to be determined beforehand by camera calibration. We used a printed chessboard pattern and the calibrateCamera function for this task. A distorted fisheye image and the result after undistortion are shown in Fig. 2(a) and Fig. 2(b), respectively. The second step transforms the undistorted image to a top-down view. This can be realized with the warpPerspective function, which takes the undistorted image and the homography matrix as input and gives the top-down view. The homography relates the undistorted and the top-down view. It can be estimated with the findHomography function and corresponding image points of the two views. We take four known corner points of the most central chessboard square with a size of (300, 300) pixels after transformation for the wanted top-down view and select four corresponding chessboard corner points from the undistorted image. The transformed image from the top-down perspective is given in Fig. 2(c).
((a)) Original image.
((b)) Undistorted.
((c)) Top-down view.
Figure 2: Transformation of the original image in a) through undistortion in b) and perspective transformation in c).
III-B QR code detection and masking
QR codes are detected using OpenCV, which outputs the four corner points of the QR code in known order. This gives the global orientation information, previously shown in Fig. 1(b). The QR code ID is extracted with the pyzbar library [^6] Availabe at: https://pypi.org/project/pyzbar/, as it gave better results than OpenCV. The QR ID enables identification of the respective chessboard square. By computing the difference in $x$ and $y$ between the square centers where the QR code is detected and where the crosshair is detected, the QR ID also gives us the square in which the crosshair lies and thus global localization capabilities.
Since the QR codes would disturb the successive image processing operations, we mask them out. Therefore, the average color around the detected QR code is computed and the region of the QR code is replaced with a slightly larger rectangle of this color. This masking procedure might also mask out parts of the crosshair if it lies on a QR code. The QR code masking could be improved in future work, but our experiments have shown that our system can cope with irregular missed crosshair detections. With some additional processing, we could also fully reconstruct a partly occluded crosshair.
III-C Crosshair detection
The green crosshair can be extracted by filtering the image with a mask that keeps the green image parts. This mask has to be carefully tuned as both the different colors of the chessboard and varying illumination condition influence the detected crosshair color. We therefore apply a concatenation of multiple masks. The first mask filters out all pixels with low green values. We then convert the image to the HSV color space, which was more robust than only working in the RBG space. Three additional masks filter out pixels with low saturation, low hue, and pixels where the sum of hue and value is low. A combined crosshair mask is found in Fig. 3(a). The horizontal and vertical lines in a region of 150 pixels around estimated crosshair position from the previous time step are then found by Hough transformation [hough_paper], clustered into lines with similar angle and start- and endpoints, and the start and end $x$ values of the clustered lines are averaged for the vertical, and the $y$ values for the horizontal line. The crosshair location in image coordinates is then estimated as the intersection of the horizontal and vertical crosshair lines. The final crosshair lines and the detected crosshair position are shown in Fig. 3(b).
((a)) Crosshair mask.
((b)) Final crosshair detection after Hough line averaging.
Figure 3: Crosshair detection process.
If the crosshair is not detected or if there is a large deviation from the expected crosshair position, we use an average of the previous crosshair positions as estimate. In principle, our system can work without the laser diode by calibration one initial crosshair and using a virtual crosshair afterwards, which stays in the same position in the image under ideal conditions. In the experiments with our low-cost robot that has limited mechanical stability, we have found that using the laser diode, which is lightweight and mounted much closer to the ground than the camera, is more accurate and robust than using a virtual crosshair due to less vibrations and shaking of the laser diode.
III-D Chessboard square detection
To detect the chessboard squares, we first apply a Canny filter after grayscale conversion and Gaussian smoothing for edge detection and mask out the previously detected laser crosshair (after extending the crosshair lines), as shown in Fig. 4(b). The binary Canny output is then processed by Hough transformation to find the chessboard lines. For further denoising, the detected lines are broadened and the image of lines is processed by an opening operation, which consists of erosion followed by dilation. In dilation and erosion, a 2D kernel is slided over the image, where the kernel center corresponds to the pixel to process. In each kernel, erosion only keeps white (detected line) pixels if all kernel pixels are white. Dilation does the opposite and sets the color of the center pixel to white if any pixel in the kernel is white. The image after Hough transformation and opening is depicted in Fig. 4(c). For further operations, the image is inverted, as seen in Fig. 4(d). The chessboard squares are detected by first finding and simplifying the white square contours with the functions findContours [find_contours] and approxPolyDP [ramer, douglas_peuker]. The simplified contours (polygons) are then filtered for square detection. Only polygons that have four edges, relative side lengths between factor 0.95 and 1.05, and an area of 80.000 - 100.000 pixels are kept. As the detected squares are slightly smaller due to the line broadening and opening, 10 pixels are added to the square corner coordinates. Those detected squares are given in Fig. 4(e). In the last step, the detected corners are refined to subpixel accuracy with cornerSubPix [subpix] shown in Fig. 4(f).
If squares are occluded, the redundancy in the chessboard pattern enables estimating all squares (including the square with crosshair used for positioning) by simply extending the detected squares to build virtual, undetected squares. Given the precise square detection, this method has proven quite accurate. If for some reason no square is detected, we apply a constant velocity model for $x$, $y$ and $\theta$ to estimate the AMR position and heading, which is sufficient for a high enough sampling frequency of camera images.
((a)) Top-down view.
((b)) Canny edges after crosshair and QR code masking.
((c)) Hough transform and opening.
((d)) Inversion.
((e)) Square detection (colored).
((f)) Corner refinement.
Figure 4: Square detection process in the perspective transformed top-down image.
III-E Position and heading estimation
Given the detected corner points of the chessboard square with detected laser crosshair, we compute the relative position of the crosshair in the square. We then convert the crosshair pixel coordinates to global coordinates with the known square size of $16.\overline{6}$ cm. To resolve the rotation ambiguity of the local square coordinate system and compute global world coordinates, the QR code references are used, as previously shown in Fig. 1(b). The heading is computed as the angle between the chessboard lines in $x$ direction and the vertical crosshair line. We currently get the chessboard lines though a line filter, but future work should simply use the chessboard corners with sub-pixel accuracy for this task.
If no QR code is detected as rotation reference, we predict the current heading using a constant velocity model. The heading is then estimated by choosing the detected angle from the set of ambiguous angles (occuring due to the lack of knowledge which chessboard lines point in $x$- and $y$-direction) by a nearest neighbor assignment. Since the AMR has to move in a physically plausible way and the heading cannot make large jumps between two images, this works reliably.
AMR position and heading estimation is also possible without QR codes through crosshair tracking in a simple fashion. Since we can reliably estimate the heading, no position ambiguity in local square coordinates occurs while tracking the crosshair as the order of P1-P4 in Fig. 1(b), i.e. the orientation of a locally observed square, stays known. To catch jumps between squares, we check if the crosshair position in local square image coordinates in $x$ or $y$ surpasses a threshold of $150$ pixels and if so, we perform a global square jump in the respective direction. [^7] A preliminary version of GG that didn’t contain QR codes yet gave similar position and heading estimation results as with QR codes, but this allows only relative position estimation w.r.t. the initial position and has no mechanism for position recovery after failure or kidnapping.
Using this crosshair tracking after an initial global position and heading estimate with a QR code, one could theoretically avoid frequent global updates with the QR code references, but this requires a high enough image sampling frequency such that the Nyquist-Shannon sampling theorem holds. The spatial frequency $f_{pattern}$ of the chessboard pattern is $f_{pattern}=\frac{1}{16.\overline{6}}\,\mathrm{cm}^{-1}$ and thus the images have to be sampled with a spatial frequency of at least $f_{image}=2f_{pattern}$ to avoid aliasing (causing jumps into the wrong square while crosshair tracking). This upper bounds the maximum AMR velocity to $v_{max}=f_{image}^{-1}f_{fps}$, where $f_{fps}$ is the sampling rate of the camera. We record with $f_{fps}=8.\overline{3}$ Hz, i.e. $v_{max}\approx 0.69\,\frac{\mathrm{m}}{\mathrm{s}}$, which is more than double of the maximum velocity of our AMR. To adapt the crosshair tracking for faster AMRs, one would have to either increase the square size, the sampling rate of the camera, or add a robust crosshair position prediction model (e.g. through odometry).
If no QR code is detected in the first frames, we assume an initial position and heading of $x,y,\theta]^{T}=[0,0,0]^{T}$ and then correct the initial estimates through rotation and translation once a QR code is available.
III-F Trajectory smoothing
To increase the accuracy of the estimated AMR positions and headings, we smooth them in post-processing. This is especially helpful in case of vibrations and shaking when the AMR moves. Since we want to use the GG system as an offline reference system for another novel, radar-based indoor localization system developed by us and described in [ipin_pascal, ipin_sven, irs25_raytracing], we choose polynomial fitting for this task. An online version, e.g. with an alpha-beta filter or with a Kalman filter [bar-shalom_estimation_2002] would, however, be straightforward.
We split the estimated trajectory into equally sized windows with 50% overlap. We then fit a polynomial of given degree to the respective $x$, $y$ and $\theta$ values in each window. We experimentally set the window size to 20 and the polynomial degree to 5. The smoothed values in the overlapping windows are averaged for better estimates and transitions between windows.
IV Experiments
IV-A Hardware
[…]
IV-B Localization experiments
To showcase the correct working of our system under real inference conditions with possible motion blur, vibrations etc., we drive the AMR in the laboratory by steering it wirelessly with a PS4 controller. Since hand-collecting ground truth on the floor [ground_truth_laser] (e.g. with a set square and a protractor) is an error prone and tedious task that only works in a static setting, we developed a software tool to label the ground truth position and heading based on the corresponding top-down image by hand.
V Results
Our trajectory results are visualized in Fig. 5.
((a)) Full crosshair trajectories.
((b)) Zoomed-in region.
Figure 5: Trajectory evaluation with a) full trajectories of the crosshair before and after smoothing and b) zoom into trajectory. Smoothing reduces the jitter in the raw trajectory caused by the shaking of the AMR.
The full true and estimated raw and smoothed trajectories are shown in Fig. 5(a). As visible, the estimates closely follow the ground truth. The raw crosshair trajectory shows some jitter. This is caused by the shaking of the low-cost robot and not by the image processing. With the proposed smoothing, the jitter is reduced and the trajectory is improved, as evident from the zoomed-in region in Fig. 5(b).
((a)) Errors.
((b)) Cumulative density function of the absolute errors.
Figure 6: Position and heading errors in a) and corresponding CDFs of absolute errors in b).
A closer error evaluation is found in Fig. 6. The errors in $x$, $y$ and $\theta$ are depicted in Fig. 6(a). All errors in $x$ and $y$ are below 1 cm. A few outliers arise due to small errors in the crosshair or corner detection, especially when their lines overlap. Heading errors are mostly below 1 degree. Heading has a few more outliers than $x$ and $y$ because the crosshair can get skewed by uneven ground or by the shaking of the AMR. Heading has also proven to be harder to label than $x$ and $y$. It should be noted that we introduce errors in the ground truth with our labeling method and we believe our system to be more accurate. We skip the error evaluation with smoothing, as it only improved the realness of the crosshair trajectory and not the accuracy. Smoothing is useful when mapping from the crosshair to a sensor from a different localization system on the AMR, where jitter degrades the accuracy of the crosshair reference. The corresponding CDFs of the absolute errors are found in Fig. 6(b). They show clearly that most errors in $x$ and $y$ are in low mm or even sub-mm range.
Table. I gives the bias, the mean absolute error (MAE), the maximum error, the 95th percentile and the 99th percentile of the absolute error for each parameter.
Table I: Error statistics for $x$ and $y$ in mm and for $\theta$ in degree.
Bias MAE 95th 99th Max
$x$ -0.02 0.89 2.96 5.04 6.44
$y$ 0.04 0.66 1.95 3.18 4.08
$\theta$ -0.35 0.51 1.25 1.65 2.11
The estimates in $x$ and $y$ have negligible bias. Positioning errors in both dimensions are similar as expected. The errors in $x$ are slightly higher due to more outliers as visible in Fig. 6(a). The 99th percentile of the absolute errors in $x$ and $y$ is far below 1 cm. The heading is underestimated as seen in the bias and the MAE is heavily influenced by this bias. Improving the laser setup, using a stable virtual crosshair, or calibration the heading by de-biasing would improve performance. Still, the 95th percentile of heading errors is just slightly over 1 and a non-negligible part of this error is likely due to the labeling process.
VI Discussion
Our proposed GG system is fast, low-cost, and gives mm accuracy. It is, however, just a prototype and this localization approach brings several advantages and can be further improved upon.
Cost and portability: Our system is low-cost due to the low hardware requirements (laser diode, camera, chessboard, processor). We are confident that the system will also work with even cheaper hardware. The largest cost factor is currently the chessboard floor. However, any grid pattern is sufficient. For rooms without such chessboard floor, one might e.g. build a grid by projection orthogonal laser lines on the floor using a laser level and drawing or taping such line pattern on the floor using the laser lines. A grid made of fiducial markers, e.g. with QR codes as in [QR_drones, QR_docking_submm], is another computationally more expensive and most likely less accurate and robust option than a grid of lines. One may also build a portable version, where pre-built chessboard grid patches can be sticked together. The system can further work with a cheaper camera (the current one already costs less than $100\,\mathdollar$) and since the image processing is relatively simple, one might implement a future (online) version that runs on a field programmable gate array (FPGA).
Accuracy and 3D position and orientation estimation: Due to the high spatial resolution of modern cameras and the closeness of the camera to the position references (chessboard corners) and the crosshair on the ground, we achieve high positioning accuracy. The localization accuracy of GG is also mostly independent of the AMR position in the room, given similar illumination conditions. Performance might be enhanced by multi-camera and multi-laser setups. The localization method could be extended to handle arbitrary detected chessboard corners, opposed to the four square corners in which the crosshair lies. Our current positioning bottleneck (with a moving AMR) is the not the localization algorithm, but the mechanical stability of the AMR, which causes larger shaking and vibration of the camera, and to a smaller extend also of the laser diode and the crosshair. Simple improvements would be to use an AMR with better stability, more rigid camera and laser diode installations, and adding some stabilizers for the camera and more importantly, for the laser diode in case of localization using a real crosshair.
An alternative localization approach that could be studied in future work is Perspective-n-Point (PnP) based localization with a monocular camera, as often used with fiducial markers. Therefore, the known correspondences between the detected 2D chessboard corner points in the image and their known 3D positions in world coordinates enable 3D camera position and orientation estimation without the laser diode.
Global localization: By equipping the floor with QR codes, we achieve global localization capabilities. QR codes were, however, just a readily available solution for this proof-of-concept of the system. Simpler visual encodings that are easier and faster to detect would improve the system, especially for an online version. The encodings can be used to either identify the chessboard squares and their corners (as in our current version), or to identify the different grid lines.
Real-time processing: In addition to an FPGA version, crosshair tracking, simpler global position and orientation references, an alpha-beta filter for AMR trajectory smoothing etc., the image processing still has much potential for improvement. Instead of running the full QR code, crosshair and corner detection for each image as described in Sec. III (which itself has room for improvement, e.g. by initial corner detection through a simpler grid pattern finding method at lower resolution instead of the expensive Hough transformation at full resolution), one might track the global references and the periodic chessboard corners after initial detection and make use of the prior knowledge about the crosshair position in the image. This would require only processing smaller patches of the images, thus speeding up the detection chain.
Robustness: The redundant, periodic chessboard pattern allows our system to work even if large areas of the chessboard pattern are occluded, e.g. by other AMRs, people, dirt etc. Since the crosshair stays roughly in the same position of the image and parts of the crosshair are sufficient to reconstruct the full crosshair, we can also deal with crosshair occlusion. Robustness could be further increased by adding more cameras, laser diodes, and odometry, which would additionally improve accuracy as noted before, but also increase system complexity and cost.
Scalability: Our GG system is scalable to large areas. In addition, the system can be scaled to robot swarms. In a version without the laser diode (virtual crosshair or PnP), there will be no interference between robots except for some floor occlusion, which the system can handle. If all robots use the same laser diode, we can still differentiate the different lasers crosshairs, as the crosshair of each respective robot stays in the same place in the images. In case there might be interference between the crosshairs, one could install lasers of different colors on the robots.
VII Conclusion
We have introduced a novel, camera-based indoor localization system called GroundGazer (GG) for 2D AMR position and heading estimation. Opposed to currently existing methods with similar mm localization and sub-degree heading accuracy, our system has much lower hardware requirements. Its limitation is the need for a chessboard or a similar grid pattern and global position and orientation references on the floor, which may limit its usability for applications where this is impractical. We built a prototype of GG using a chessboard floor with known corner coordinates as passive position references, QR codes as global position and orientation references, a ground gazing fisheye camera that is placed on the AMR at low height, and a laser crosshair that is projected from the AMR onto the chessboard floor and observed by the camera for position and heading estimation. The prototype achieved low mm localization and sub-degree heading estimation accuracy.
Our prototype will be used in future work to benchmark our radar-based indoor localization system. We have further discussed the many advantages this localization approach brings and how it may be improved and extended in various directions in future work.