### v5

# GroundGazer: An optical reference system for planar localization with millimeter accuracy at low cost

Abstract
Highly accurate indoor localization systems with absolute mm positioning accuracy are currently expensive. They include laser trackers, total stations, and motion capture systems relying on multiple high-end cameras. In this work, we introduce a high-accuracy, planar indoor localization system named GroundGazer (GG) for autonomous mobile robots (AMRs). GG estimates the AMR’s planar position with mm and its heading with sub-degree accuracy. The system requires only a monocular (fisheye) camera, a chessboard floor, and an optional laser diode. Our system is simple and low-cost due to the chessboard floor, robust, scalable to multiple robots, and extendable to 3D position and orientation estimation.
Keywords— Indoor localization, mobile robots, computer vision
I Introduction
Benchmarking lower-accuracy indoor positioning systems with up to cm localization accuracy requires localization approaches with reliable mm accuracy. Currently, such systems (described in Sec. II-B) are very expensive. Costs for deploying a motion capture system, a common solution for accurate position referencing, can easily exceed $\mathdollar 10000$, even for a small laboratory. While some alternatives might be cheaper, they usually come with other drawbacks such as strict line of sight (LOS) requirements, or degradation with multi-path.
I-A Contributions
In this work, we introduce a camera-based, planar indoor localization system for AMRs which achieves mm localization and sub-degree heading accuracy at multitudes lower cost than existing systems of similar precision for small rooms. We name it GroundGazer, as the camera placed on the AMR gazes at the passive position references on the ground. Its intended application is to serve as reference for less accurate localization methods. The proposed GG prototype requires only a chessboard floor (or similar grid pattern), a camera (here a global shutter RGB fisheye camera), and an optional laser diode, i.e. low-cost, off-the-shelve components. The current implementation estimates planar position ($x$, $y$) and heading $\theta$, which is sufficient for planar AMR localization with known or irrelevant height, roll, and pitch. In Sec. VI, we discuss an extension to 3D position and orientation estimation, at the cost of higher computational complexity and possibly reduced accuracy.
The idea for GG originates from [2]. To evaluate the accuracy of their proposed positioning system, they installed two point laser diodes that project two points from the front and the back of their localized vehicle onto the chessboard floor beneath it. The manually collected positions of those two points w.r.t. the known corner coordinates of the chessboard enable precise position and heading referencing. Our GG system relies on the same idea, resulting in a highly accurate, automated, camera-based indoor localization system.
I-B System overview
The GG prototype is depicted in Fig. 1(a).
((a)) Our system in the laboratory equipped with a chessboard floor and QR codes. The AMR has a ground gazing camera and a laser diode that projects a crosshair on the floor at the front.
((b)) 2D position and heading estimation using the detected crosshair (P5), the chessboard square corners (P1-P4), and the $x$-direction chessboard lines. The globally known square corner coordinates through the QR codes enable accurate, global positioning.
Figure 1: Overview of our GG localization system in a) and conceptual depiction of the localization method in b).
[…]
The localization system works by detecting the crosshair in a chessboard square. We selected a blue and white chessboard floor and a green laser diode, as green lasers are most easily visible and other commercially available black and white chessboard floors complicate the crosshair detection, caused by the large difference in laser reflective behavior between black and white squares. The camera image is calibrated such that the orientation of the chessboard in the image directly gives the heading. The corner coordinates of the chessboard squares are known and serve as passive position references for AMR localization.
Since the camera only observes a small, local area of the periodic chessboard pattern, the corner coordinates are ambiguous. Adding different QR codes as global position and orientation references to the floor resolves this ambiguity. All QR codes are coarsely aligned with the chessboard lines, and their corners can thus be used to extract the $x$- and $y$-directions of the chessboard lines in global coordinates. The QR code IDs allow for identification of the QR code squares and global positioning. By measuring the offset between a detected QR code square and the square containing the crosshair, the crosshair square and its known corner coordinates can be identified without requiring a QR code in each square.
The AMR position is defined as the crosshair center location P5. Its planar position $(x,y)$ is given by the relative position of the crosshair w.r.t. its square corners P1-P4 and the heading $\theta$ is given by the angle between the chessboard lines in $x$-direction and the vertical image axis. An image describing the AMR positioning is shown in Fig. 1(b).
GG’s system design enables high accuracy with simple visual infrastructure. While the requirement of a chessboard floor limits GG’s use as a general localization method, deploying such chessboard floor comes at reasonable effort and at low cost for small areas. GG is not restricted to visible light and could also be realized with an active infrared (IR) camera and a passive IR grid, improving robustness to ambient illumination and allowing the positioning infrastructure to remain largely unobtrusive for human users. The image processing to estimate the AMR’s position and heading is described in Sec. III.
II Related works
We focus on closely related systems in this section. For an overview of general indoor localization technologies, we refer to the surveys in [14] and [21], and for optical indoor localization systems to [12] and [15].
II-A Related works with ground gazing cameras
Infrastructure-free approaches localize using only the given ground texture. Relative positioning can be realized through simultaneous localization and mapping (SLAM) or feature-matching in successive images to build a vision-based odometer [17]. Absolute positioning is possible with feature-matching of images collected during inference with a pre-collected image database of the ground [23, 4, 20]. However, mapping a floor and anchoring it to the real world with correct absolute scale requires an expensive measurement campaign with accurate ground truth, e.g. from motion capture systems [4].
While feature-matching requires no infrastructure, using known position references is the more suitable approach for reference systems due to higher accuracy, robustness, easier setup, and lower hardware requirements. Various systems place infrastructure (fiducial markers) on the floor for absolute, global localization of AMRs [9], and other robots, such as walking assistants [16] and drones [8]. By applying popular markers including QR codes, ArUco markers, or AprilTags, they achieve dm to cm localization accuracy. A QR code-based docking system in [3] reports sub-mm accuracy with very small camera-QR code distance. Although low-cost, the very small camera-QR code distance and the need for many QR codes limit its use for AMR localization. Visual floor marker grids for AMR navigation can also be found in commercial systems, e.g. the Hercules AMR by Amazon robotics.
The chessboard line grid in GG is easier to realize than a precisely placed marker grid and offers more accurate corner detection with high robustness against occlusion due to its redundant pattern. In a laboratory, it can also facilitate installation of other equipment.
II-B Localization technologies with mm positioning accuracy
Optical systems include motion capture systems like the commercial systems by Qualisys and Vicon [^1] Available at: https://www.qualisys.com/ and https://www.vicon.com/.. They consist of passive IR markers placed on the tracked AMR. With one marker, 3D positioning is possible. A minimum of three non-collinear markers enable 3D position and orientation estimation. While capable of sub-mm 3D localization and sub-degree orientation estimation, they require LOS, multiple high-end IR cameras, and become very expensive for large-scale deployments. Further, since the cameras are not installed around the AMR, accurate time synchronization is required to ensure mm accuracy. A different optical solution are fiducial markers. In [7], sub-mm accuracy is achieved, but only with costly cameras and metrology-grade equipment to precisely install custom reference markers. The previously mentioned QR code docking system in [3] also gives sub-mm accuracy, but is limited by camera-QR code distance.
Total stations (tachymeters) pose another solution for 3D position estimation with mm or sub-mm accuracy, e.g. as used in [5]. Total stations enable long range sensing, but like motion capture systems, they are expensive and require LOS and accurate synchronization.
In acoustic ultrasound systems, the low propagation speed of sound greatly reduces sensitivity to timing and synchronization errors compared to radio frequency (RF)-based ultra wideband (UWB) systems. Using time of flight (ToF)-based measurements between multiple active anchors and the tracked user, a commercial system reports real-time 3D positioning accuracy of up to 1.5 mm using 6 acoustic anchors with ideal geometric placement [22].
While ToF-based localization with UWB systems only gives dm to cm accuracy, recent studies show mm accuracy by combining initial, coarse ToF-based range estimates with refined, ambiguity-resolved phased-based range estimates [11]. Other RF-based systems include the interferometry-based system developed by Koherent [^2] Available at: https://koherent.io/, and frequency-modulated continuous wave (FMCW) radar-based backscatter systems using reflective tags that modulate the reflected signal [1, 18]. Deep learning-based channel state information (CSI) fingerprinting with a massive multiple input multiple output (MIMO) system could recently also achieve mm accuracy in a small area and under LOS conditions [10]. However, domain shifts between training and inference, caused by environmental changes, reduce the accuracy of such systems. RF-based systems often require specialized hardware and degrade under multipath. At smaller wavelength, using light detection and ranging (LiDAR) with reflectors at known positions as references has also shown mm accuracy [19], but LiDAR has high sensor cost.
Additional systems with (sub-)$m$ accuracy based on visual markers, laser projections, magnetic fields, laser trackers, pseudolites, and infrared laser beams (iGPS) can be found in [12] and [13].
III GroundGazer
III-A Image transformation
The initial step of our GG localization algorithm transforms the distorted and skewed image collected by the fisheye camera to a top-down view of the chessboard floor. This step consists of two transformations. First, the fisheye image is undistorted. The required intrinsic matrix $\mathbf{K}$ of the camera and its distortion coefficient vector $\underline{d}$ are estimated with a printed chessboard pattern. A distorted fisheye image and the undistorted result are shown in Fig. 2(a) and Fig. 2(b), respectively. Then, the undistorted image is transformed to a top-down view by perspective transformation, depicted in Fig. 2(c). Since GG localizes the crosshair center relative to its square corners, errors in undistortion and perspective transformation have little influence on the positioning accuracy. The homography matrix for perspective transformation is estimated offline with four corresponding square corner points of both views. The AMR heading has to match the orientation of the transformed chessboard to align them. We calibrate at $\theta=90^{\circ}$, thus the edges of the top-down view square are parallel to the image axes. We set its size to (300, 300) pixels. The four corresponding points from the undistorted image view are the corners of the central crosshair square.
Alternatively, heading could be estimated with the angle between a similarly calibrated, vertical crosshair line that points in heading direction and the $x$-direction chessboard line. This makes heading estimation independent of the camera orientation and thus of camera vibrations. However, it is still susceptible to lesser laser diode vibrations and it makes heading estimation dependent on the crosshair line detection, which is worse than the chessboard line detection due to broader edges, crosshair distortion on uneven ground etc., and therefore not chosen.
((a)) Original image.
((b)) Undistorted.
((c)) Top-down view.
Figure 2: Transformation of an original image in a) through undistortion in b) and perspective transformation in c).
III-B QR code detection and masking
QR code detection gives the four corner points of a QR code in known order and its ID, which enables identification of all chessboard squares and their corners.
As the QR codes would disturb successive image processing, we mask them out. Our procedure might also mask parts of the crosshair if it lies on a QR code and could be improved in future work. Nevertheless, our experiments have shown that GG can cope with irregular missed crosshair detections and by proper parametrization of the Hough transform in the crosshair detection, we can also detect an incomplete crosshair.
III-C Crosshair detection
The green crosshair can be extracted by filtering the image with a mask that keeps the green image parts. This mask has to be carefully tuned, as both the different colors of the chessboard and varying illumination condition influence the detected crosshair color. We therefore apply a concatenation of multiple masks. The first mask filters out all pixels with low green values. We then convert the image to the HSV color space, which was more robust than only working in the RBG space. Three additional masks filter out pixels with low saturation, low hue, and pixels where the sum of hue and value is low. A combined crosshair mask is found in Fig. 3(a). The horizontal and vertical lines in a region of 150 pixels around the estimated crosshair position from the previous time step are then found by probabilistic Hough transformation, clustered into lines with similar angle and start- and endpoints, and the start and end $x$-values of the clustered lines are averaged for the vertical, and the $y$-values for the horizontal line. The crosshair location in image coordinates is then estimated as the intersection of the horizontal and vertical crosshair lines. The crosshair lines and the detected crosshair position are shown in Fig. 3(b).
((a)) Crosshair mask.
((b)) Final crosshair detection.
Figure 3: Crosshair detection.
If the crosshair is not detected or at an unexpected position, we use an average of the previous crosshair positions as estimate. In principle, our prototype can work without the laser diode by replacing the real crosshair with a virtual one, which stays in the same position in the image under ideal conditions (i.e. with a perfectly stabilized camera). This means selecting one point in image coordinates to replace the crosshair center, and using it as AMR position. Heading is estimated based on the chessboard rather than the crosshair orientation. On our low-cost robot, the lower mounted laser diode proved to be more accurate and robust than a virtual crosshair, being less affected by vibrations under AMR motion.
III-D Chessboard square detection
To detect the chessboard squares, we first apply a Canny filter for edge detection after grayscale conversion and Gaussian smoothing. The binary Canny output is then processed by probabilistic Hough transformation to find the chessboard lines. For further denoising, the detected lines are broadened and the image of lines is processed by an opening operation. The image after Hough transformation and opening is depicted in Fig. 4(b). We then invert the image. The chessboard squares are detected by first finding and simplifying the white square contours. The simplified contours (polygons) are then filtered for square detection, where only polygons with four edges, relative side lengths between factor 0.95 and 1.05, and an area of 80.000 - 100.000 pixels are kept. Finally, the detected corners are refined to sub-pixel accuracy [6], shown in Fig. 4(c).
If squares are occluded, the redundancy in the chessboard pattern enables estimating all squares by simply extending the detected squares to build virtual, undetected squares. Using the precise square detection, this method has proven quite accurate. If for some reason no square is detected, we apply a constant velocity model for $x$, $y$ and $\theta$, which is sufficient with a high enough sampling frequency of camera images.
((a)) Top-down view after QR code masking.
((b)) Hough lines after opening.
((c)) Square detection (red).
Figure 4: Square detection in the top-down image.
III-E Position and heading estimation
Given the detected corner points of the chessboard square with detected laser crosshair, we compute the relative position of the crosshair in the square. We then convert the crosshair pixel coordinates to global coordinates with the known square size of $16.\overline{6}$ cm. To resolve the rotation ambiguity of the local square coordinate system and compute global world coordinates, the QR code references are used, as previously shown in Fig. 1(b). The heading is computed as the angle between the chessboard lines in $x$-direction and the calibrated vertical image direction.
If no QR code is detected, we predict the current heading using a constant velocity model. The heading is then estimated by choosing the detected angle from the set of four ambiguous angles (due to the lack of knowledge which chessboard lines point in $x$- and $y$-direction) by a nearest neighbor assignment. The four ambiguous chessboard line directions are computed by clustering and averaging the lines from Hough transformation in Fig. 4(b). Since the AMR has to move in a physically plausible way and the heading cannot make large jumps between two images, this nearest neighbor assignment works reliably.
AMR position estimation is also possible without QR codes through crosshair tracking in a simple fashion. Since we can reliably estimate the heading, no position ambiguity in local square coordinates occurs while tracking the crosshair, as the order of P1-P4 in Fig. 1(b), i.e. the orientation of a locally observed square, stays known. To catch jumps between squares, we check if the crosshair position in local square image coordinates in $x$ or $y$ surpasses a threshold of $150$ pixels and if so, we perform a global square jump in the respective direction.
Using this crosshair tracking after an initial global position and heading estimate with a QR code, further frequent global updates with the QR code references are not necessary, but this requires a high enough image sampling frequency such that the Nyquist-Shannon sampling theorem holds. The spatial frequency $f_{pattern}$ of the chessboard pattern is $f_{pattern}=\frac{1}{16.\overline{6}}\,\mathrm{cm}^{-1}$ and thus the images have to be sampled with a spatial frequency of at least $f_{image}=2f_{pattern}$ to avoid aliasing (causing jumps into the wrong square during crosshair tracking). This upper bounds the maximum AMR velocity to $v_{max}=f_{image}^{-1}f_{fps}$, where $f_{fps}$ is the sampling rate of the camera. We record with $f_{fps}=8.\overline{3}$ Hz, i.e. $v_{max}\approx 0.69\,\frac{\mathrm{m}}{\mathrm{s}}$, which is more than double of the maximum velocity of our AMR. To adapt the crosshair tracking for faster AMRs, one would have to either increase the unambiguous region (e.g. by increasing the square size), increase the sampling rate of the camera, or integrate an accurate kinematic model (e.g. through odometry).
If no QR code is detected in the first frames, we assume an initial position and heading of $x,y,\theta]^{T}=[0,0,0]^{T}$ and then correct the initial estimates through rotation and translation once a QR code becomes available.
IV Experiments
IV-A Hardware
Our prototype AMR is a low-cost Raspberry Pi robot (SunFounder Picar-X, $\sim\mathdollar 100$) with Ackermann kinematics. The USB RGB fisheye camera is from Svpro $(\sim\mathdollar 100)$ and offers a resolution of 1920x1200 pixels, a FoV of $150^{\circ}$, and fast image creation (i.e. negligible motion blur) by employing a global shutter. The green laser diode is from CTRICALVER $(<\mathdollar 10)$ with a wavelength of $520\text{\,}\mathrm{n}\mathrm{m}$ and has $1\text{\,}\mathrm{m}\mathrm{W}$ of power. The chessboard floor is made from Polyvinylchlorid (PVC) with a known square size of ($\overline{6}, 16.6¯1$$\overline{6}$) cm.
IV-B Localization experiments
To showcase the correct working of our system under real inference conditions with possible motion blur, AMR vibrations etc., we drive the AMR in the laboratory by steering it wirelessly with a PS4 controller. Since we have no access to a costly high-accuracy reference system, which was our motivation to design the low-cost and high-accuracy GG, we developed a software tool to label the ground truth position and heading based on the corresponding top-down image by hand, similar to the annotation procedure of various localization approaches based on pixel differences [23, 20]. We therefore zoom into the wanted points and select the crosshair center position, and three corners of the respective square to build the coordinate system for localization, including the $x$-direction line for heading estimation. The points are selected to the best of our abilities with pixel accuracy, where one pixel equals roughly $\frac{16.66}{300}\approx 0.56$ mm (see Fig. 1(b)). Future work should benchmark GG against external ground truth, potentially in a version more robust to AMR vibrations discussed in Sec. VI.
V Results
The labeled and estimated trajectories are shown in Fig. 5. The estimates closely follow our hand-labeled trajectory.
Figure 5: Trajectory evaluation.
((a)) Errors.
((b)) Cumulative distribution functions of the absolute errors.
Figure 6: Position and heading errors in a) and corresponding CDFs of absolute errors in b).
A closer error evaluation is found in Fig. 6. The errors in $x$, $y$ and $\theta$ are depicted in Fig. 6(a). A few outliers arise due to small errors in the crosshair or corner detection, often caused by overlaps of the crosshair with the lines and corners of the chessboard. The corresponding CDFs of the absolute errors are found in Fig. 6(b). They show clearly that most errors in $x$ and $y$ are in low mm or even sub-mm range.
Table. I gives the bias, the mean absolute error (MAE), the maximum absolute error (Max), and the 95th and 99th percentiles of the absolute error for each parameter.
Table I: Error statistics for $x$ and $y$ in mm and for $\theta$ in degrees.
Bias MAE 95th 99th Max
$x$ -0.03 0.92 2.62 5.03 6.44
$y$ 0.02 0.69 1.89 3.02 3.70
$\theta$ -0.05 0.36 0.87 1.30 1.39
All estimates have negligible bias. Positioning errors in both dimensions are similar as expected. The errors in $x$ are higher due to more outliers as visible in Fig. 6(a). The 99th percentile of the absolute errors and the max errors in $x$ and $y$ are well below 1 cm. Heading estimation also shows good performance with the 95th percentile of absolute heading errors being slightly under $1^{\circ}$. Good heading accuracy is important to accurately map from the crosshair center to another position around the crosshair, e.g. to a localization sensor on the AMR.
VI Discussion
Our proposed GG system is low-cost and gives mm accuracy. It is, however, just a prototype and this localization approach brings several advantages and can be further improved upon.
Cost and portability: The largest cost factor of GG is the chessboard floor, which is still cheaper than using other mm-accuracy methods for small rooms. Rooms without a chessboard floor could instead use a taped or drawn grid of lines. By connecting pre-built chessboard patches, one could also port the system.
Accuracy: Positioning accuracy is mostly independent of the AMR position in the room, given similar illumination conditions. Current localization requires four crosshair square corners and might be extended to accept any (and less) corners. This allows crosshair projection closer to the AMR, or replacing the crosshair center with a known, visible point on the AMR (see e.g. Fig. 2(a) bottom) for position referencing. The error in mapping from GG’s position to a sensor on the AMR could also be minimized with a camera mount beneath the AMR. However, this reduces FoV and the feasible chessboard square size. It also increases motion blur, necessitating reduced exposure time and thus additional lighting or camera flash. The current limitation during motion is the mechanical stability of the AMR, causing camera and crosshair jitter. Since we only evaluate the crosshair position accuracy, not a jittered projection to a stable position on the AMR, this effect is currently not measured. Such jitter could be reduced with more stable AMRs, rigid camera and laser diode mounts, mechanical stabilizers, and by smoothing the trajectory. Replacing the crosshair center with a position reference on the AMR, or with a stabilized point in image coordinates, makes the laser diode obsolete.
3D positioning: Future work could study Perspective-n-Point (PnP)-based localization, as often used with fiducial markers. The known camera intrinsics and correspondences between the detected 2D image chessboard corner points and their known 3D world positions enable 3D camera position and orientation estimation. PnP requires no laser diode and might also reduce the error from projecting the estimated position by GG to the position of another localization sensor on the AMR.
Global localization: By placing QR codes on the floor, GG achieves absolute, global localization. QR codes were, however, just a convenient solution for this proof-of-concept. Simpler visual encodings with faster detection would improve the system, especially for online operation. The encodings serve to identify the chessboard square corners, e.g. with the known QR corners as in our current version, or by alternatives such as identifiable grid lines.
Real-time processing: Instead of running the full detection chain for each image as described in Sec. III, one could track the global references and chessboard corners after initial detection. This requires processing only smaller image patches, thus speeding up the detection chain.
Robustness: The redundant, periodic chessboard pattern allows GG to operate even when large areas of the chessboard are occluded, e.g. by other AMRs, people, dirt etc. Robustness could be increased by adding more cameras, laser diodes, and odometry, which would additionally improve accuracy, but also increase system complexity and cost.
VII Conclusion
We have introduced a camera-based indoor localization system called GroundGazer (GG) for 2D AMR position and heading estimation. Opposed to currently existing methods with similar mm localization and sub-degree heading accuracy, our system has much lower hardware requirements. Its limitation is the need for a chessboard or a similar grid pattern and global position and orientation references on the floor, which may limit its usability for applications where this is impractical. We built a prototype of GG using a chessboard floor with known corner coordinates as passive position references, QR codes as global position and orientation references, a ground gazing fisheye camera placed at low height on the AMR, and a laser diode that projects a crosshair from the AMR onto the chessboard floor. The crosshair and the floor are continuously observed by the camera for position and heading estimation. Our prototype achieved low mm localization and sub-degree heading accuracy. We have further discussed the advantages this localization approach brings and how it may be improved and extended in various directions in future work.